OpenStudy (rainbow_dash):

Laura has moved to a new apartment. Her schoolbooks comprising of different subjects are mixed in a bag during the move. Four books are of mathematics, three are English, and six are science. If Laura opens the bag and selects books at random, find the given probability. P(3 math books)

OpenStudy (anonymous):

Since there are two different answers, I will see if I can get the right answer. P(1 science and 2 mathematics books) = probability of getting a math book * p(second math book) * p(science book)? This is the solution to one of the answers I see on here. Although it seems correct, its missing one very important thing. What if you didn't get you a math book first but a science book instead. So the correct solution is 4/13 * 3/12 * 6/11 * 3 = 216/1716 = 36/286 I multiplied by 3 because you can have 3 different possibilities: math book math book science book science book math book math book math book science book math book Hope you understand this a bit better!